FTMW SPECTROSCOPY OF CH_{2}NC (X̃^{2}B_{1}) ∼ A POSSIBLE CANDIDATE OF INTERSTELLAR MOLECULE

\underline{T. HIRAO}, Institute for Astrophysics and Planetary Sciences, Ibaraki University, 2-1-1 Bunkyo, Mito, Ibaraki 310-8512, Japan.; H. OZEKI, Japan Aerospace Exploration Agency (JAXA), SMILES Mission Team, Tsukuba Space Center, 2-1-1 Sengen, Tsukuba 305-8505 Japan.; S. YAMAMOTO, Department of Physics and Research Center for the Early Universe, The University of Tokyo, Bunkyoku, Tokyo 113-0033, Japan.

Cyanohydrocarbons are molecules of great importance in cold dark clouds such as TMC-1. Due to such interests, CH_{2}CN has been studied by experimental and theoretical means, and these extensive investigations lead unambiguous identification of the molecule in interstellar matter^{a}. It turns out that an isomer CH_{2}NC was rarely studied experimentally, probably because of unstableness of this species compared with CH_{2}CN.

We recently recorded Fourier transform microwave (FTMW) spectra of CH_{2}NC at the University of Tokyo. About 0.3∼1 percent of CH_{3}NC, synthesized by an ordinary way, was diluted in the Ar or Ne buffer gas at the pressure of ∼1 atm. The mixed gas was injected into the Fabry-Pérot cavity via a pulsed nozzle, whose opening was synchronized with high voltage dc pulse to discharge to generate the aimed species. N=1_{10}-0_{00} (∼ 22.2 GHz), 2_{02}-1_{01}, 2_{12}-1_{11}, and 2_{11}-1_{10} (∼ 44.4 GHz) were recorded with well resolved fine and hyperfine structure. The molecular structure as well as a possibility of detection of this molecule will be discussed.

aW. M. Irvine et al. Astrophys. J., 334, L107 (1988).